# Weight space Detection of Backdoors in LoRA Adapters

### v1

@@ 1 Introduction
LoRA adapters Hu et al. (2021) allow efficient fine-tuning of large language models and are widely shared through platforms like Hugging Face Hub. However, this creates a security risk: attackers can upload poisoned adapters that behave normally until a specific trigger appears in the input (Gu et al., 2019; Kurita et al., 2020). For example, an adapter might output “HACKED” whenever it sees the token “cf”. Since only the small adapter is modified while the base model stays frozen, these backdoors are hard to detect through manual inspection. A single poisoned adapter downloaded by thousands of users could compromise many downstream applications.
@@
We propose a detection method that analyzes LoRA weight matrices directly, without running the model. Our key insight is that backdoors leave a distinctive pattern in the weights: the singular values become concentrated (high energy in few values, low entropy) (Tran et al., 2018). This happens because backdoor tasks are simple mappings (trigger $\rightarrow$ response) that dominate the weight update (Luong and Chen, 2026). We extract five statistics from the weight matrix and flag adapters that deviate from normal patterns.
@@ 2 Related Work
Backdoor defenses. Existing defenses fall into three categories, all unsuitable for hub-scale LoRA screening: (1) data-centric methods that filter training data; (2) activation-based monitors requiring model execution Sun et al. (2025); and (3) weight-space methods that either need a clean reference model Chen et al. (2018) or perform expensive trigger inversion. None offer a static, data-agnostic analysis of adapter weights.
Spectral signatures and weight-space detection. The core hypothesis of our method—that backdoors leave a spectral trace—is grounded in prior work. Tran et al. (2018) showed poisoned data creates outliers in feature covariance. Recent analysis by Luong and Chen (2026) provides a spectral explanation for LoRA’s vulnerability, showing backdoor tasks create distinct geometric signatures in the weight matrices. We leverage this principle to build the first fully static detector that analyzes LoRA weights ($\Delta W$) without model execution.
@@ 3 Method
Given a LoRA adapter, we compute the weight update $\Delta W=BA$ where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$. For Llama-3.2-3B, the hidden dimension $d=4096$ and we use rank $r=16$. We extract $A$ and $B$ from the adapter’s safetensors file and compute $\Delta W$ for each attention projection (query, key, value, output), then sum them:

### v2

@@ 1 Introduction
LoRA adapters Hu et al. (2021) allow efficient fine-tuning of large language models and are widely shared through platforms like Hugging Face Hub. However, this creates a security risk: attackers can upload poisoned adapters that behave normally until a specific trigger appears in the input (Gu et al., 2019; Kurita et al., 2020). For example, an adapter might output “HACKED” whenever it sees the token “cf”. Since only the small adapter is modified while the base model stays frozen, these backdoors are hard to detect through manual inspection. A single poisoned adapter downloaded by thousands of users could compromise many downstream applications. The risk of poisoned adapters is further compounded by recent findings that open-weight LLMs can exhibit evaluation-aware behavior that scales predictably with model size Chaudhary et al. (2025), suggesting that malicious adapters may be capable of concealing backdoor behavior during evaluation while remaining active in deployment—making static, pre-deployment weight-space screening all the more critical.
@@
We propose a detection method that analyzes LoRA weight matrices directly, without running the model. Our key insight is that backdoors leave a distinctive pattern in the weights: the singular values become concentrated (high energy in few values, low entropy) (Tran et al., 2018). This happens because backdoor tasks are simple mappings (trigger $\rightarrow$ response) that dominate the weight update (Luong & Chen, 2026). We extract five statistics from the weight matrix and flag adapters that deviate from normal patterns.
@@ 2 Related Work
Backdoor defenses. Existing defenses fall into three categories, all unsuitable for hub-scale LoRA screening: (1) data-centric methods that filter training data; (2) activation-based monitors requiring model execution Sun et al. (2025); Chaudhary & Barez (2025); and (3) weight-space methods that either need a clean reference model Chen et al. (2018) or perform expensive trigger inversion. None offer a static, data-agnostic analysis of adapter weights.
Spectral signatures and weight-space detection. The core hypothesis of our method—that backdoors leave a spectral trace—is grounded in prior work. Tran et al. (2018) showed poisoned data creates outliers in feature covariance. Recent analysis by Luong & Chen (2026) provides a spectral explanation for LoRA’s vulnerability, showing backdoor tasks create distinct geometric signatures in the weight matrices. We leverage this principle to build the first fully static detector that analyzes LoRA weights ($\Delta W$) without model execution.
@@ 3 Method
Given a LoRA adapter, we compute the weight update $\Delta W=BA$ where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$. For Llama-3.2-3B, the hidden dimension $d=4096$ and we use rank $r=16$. We extract AA and $B$ from the adapter’s safetensors file and compute $\Delta W$ for each attention projection (query, key, value, output), then sum them: